# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ A-E Maximum A Posteriori (MAP)
Input: $\mathbf{X},Y,\mathbf{C},\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N},\left\{\mathbf{\Phi}_{i}\right\}_{i=1}^{N}$
@@
Appendix B Probability of Error (Proof of Lemma 1)
@@
Let $i\in\left[K\right],j\in\left[\binom{N-K}{i}\right]$. Let ~$\tilde{\mathcal{I}}^{j}_{dif}$ denote the set of all indices $\omega$ that result in a the set of FN of size $i$ as in $j$, i.e.
~$\tilde{\mathcal{I}}^{j}_{dif}=\left\{\omega\in\left[\binom{N}{K}\right]:\quad FN\left(\omega,\omega^{*}\right)=\Omega_{i}\left(j\right)\right\}.$ (15)
@@
Recall $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. To visualize this definition, consider the example in Section IV-A. For $S_{\omega^{\star}}=\left\{1,2,3\right\}$, let $j$ be the index such that $\Omega_{2}\left(j\right)=\left\{2,3\right\}$. In this case, ~$\tilde{\mathcal{I}}^{j}_{dif}$ has all indices that point to $\left\{1,4,5\right\},\left\{1,4,6\right\},\left\{1,4,7\right\},\left\{1,5,6\right\},\left\{1,5,7\right\},\left\{1,6,7\right\}$.
@@
By substituting Eq. (17) in Eq. (B), it follows that
$P\left(E_{i}^{j}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$ (18)
$\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
$=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
$\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
$\qquad\forall s>0,\rho\in\left[0,1\right]$
@@
$P\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ (19)
$=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
$\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
$=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
@@
$\begin{split}&\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\},\\ &\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\},\\ &\text{and}\\ &\omega\in\left[\binom{N}{K}\right]\;s.t.\;S_{\omega}=S_{1}\cup S_{2}.\end{split}$
@@
$\binom{N}{K-i}\cdot\binom{N-\left(K-i\right)}{i}$

### v5

@@ A-F Maximum A Posteriori (MAP)
Input: $\mathbf{X},Y,\mathbf{C},{\color[rgb]{0,0,0}\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}}$
@@
Appendix B Proof of Theorem 3
Let $M_{0}$ denote the number of negative tests, and let $G$ denote the number of non-defective items that never appear in any negative test. We follow the $(M_{0},G)$ conditioning framework for DD given in [3, Appendix A.C]. Throughout the proof, we condition on a fixed realization of the number of defectives $k$; the Bernoulli design parameter is $p=\ln(2)/K_{\max}$, but the argument below holds for any $p\in(0,1)$. For a fixed defective item $i$, let $L_{i}$ be the number of positive tests that: (i) contain item $i$, and (ii) contain no other item from $\mathcal{P}^{(S_{1,1})}$.
We condition on $(M_{0},G)=(m_{0},g)$ and analyze a fixed defective item $i$. Let $\mathsf{Pos}$ denote the event that a given test is positive, and let $\mathsf{SoloDef}(i)$ denote the event that the test contains item $i$ and no other defective items. Under a Bernoulli(p)(p) design with $k$ defectives, $q_{0}\triangleq\Pr(\text{test negative})=(1-p)^{{\color[rgb]{0,0,0}k}}$ denotes the probability that a test is negative, and $q_{1}\triangleq\Pr(\mathsf{SoloDef}(i))=p(1-p)^{{\color[rgb]{0,0,0}k}-1}$ denotes the probability that a test contains defective $i$ and no other defectives, so that
$a\triangleq\Pr(\mathsf{SoloDef}(i)\mid\mathsf{Pos})=\frac{\Pr(\mathsf{SoloDef}(i))}{\Pr(\mathsf{Pos})}=\frac{q_{1}}{1-q_{0}},$
is the effective probability that a test includes item $i$ and no other defective items, given that the test is positive. Given $G=g$, the probability that none of the $g$ surviving non-defective candidates is included in a test is $(1-p)^{g}$. Hence, for each positive test, the probability that it contains item $i$ and no other item from the Stage 1 candidate set equals $a(1-p)^{g}$. Since there are $T-m_{0}$ positive tests and the Bernoulli design is independent across tests, it follows that
$L_{i}\,\big|\,(M_{0}=m_{0},G=g)\sim\mathrm{Bin}\!\big(T-m_{0},\ a(1-p)^{g}\big).$
Thus, we have
$\color[rgb]{0,0,0}P(L_{i}>0\mid M_{0}=m_{0},G=g)=1-P(L_{i}=0\mid M_{0}=m_{0},G=g).$}
For a binomial random variable, $P(L_{i}=0)=(1-\theta)^{n}$ with $n=T-m_{0}$ and $\theta=a(1-p)^{g}$, as detailed in Appendix B-A, we obtain
$P(L_{i}>0\mid M_{0}=m_{0},G=g)\;=\;1-\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}}.$ (18)
By symmetry of the items, the expected number of items identified by the DD step is
$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]\;=\;{\color[rgb]{0,0,0}k}\,P(L_{1}>0).$ (19)
We now apply Lemma 2, which states that for any $u\in[0,1]$ and integer $n\geq 1$
$1-(1-u)^{n}\geq nu-\frac{n(n-1)}{2}u^{2}.$
Applying this with $u=a(1-p)^{g}$ and $n=T-m_{0}$ to (18) yields
$P(L_{1}>0\mid M_{0}=m_{0},G=g)$ $\geq(T-m_{0})\,a(1-p)^{g}-\frac{(T-m_{0})(T-m_{0}-1)}{2}\,a^{2}(1-p)^{2g}.$ (20)
Taking expectation of both sides of (20) with respect to $(M_{0},G)$, substituting into (19), and by linearity, we obtain
$\mathbb{E}\big[\,|\mathcal{P}^{(\mathrm{DD})}|\,\big|\,K={\color[rgb]{0,0,0}k}\big]$ $={\color[rgb]{0,0,0}k}\,\mathbb{E}\big[P(L_{1}>0\mid M_{0},G)\big]$
$\geq{\color[rgb]{0,0,0}k}\left\{\,\mathbb{E}\big[(T-M_{0})a(1-p)^{G}\big]-\,\mathbb{E}\left[\frac{(T-M_{0})(T-M_{0}-1)}{2}\,a^{2}(1-p)^{2G}\right]\right\}$
$={\color[rgb]{0,0,0}k}\left\{a\,\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]-\frac{a^{2}}{2}\,\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]\right\}.$ (21)
Conditioned on $M_{0}=m_{0}$, each non-defective item appears in none of the $m_{0}$ negative tests with probability $(1-p)^{m_{0}}$. Hence, the number $G$ of such non-defectives follows
$G\mid M_{0}=m_{0}\;\sim\;\mathrm{Bin}(N-{\color[rgb]{0,0,0}k},\,(1-p)^{m_{0}}).$
Let $\rho=(1-p)^{m_{0}}$ denote this probability. Then, the probability generating function of a binomial variable implies
$\mathbb{E}\big[(1-p)^{G}\mid M_{0}=m_{0}\big]=\mathbb{E}\big[(1-p)^{G}\big]=\left(1-p\cdot\rho\right)^{N-{\color[rgb]{0,0,0}k}}=\left(1-p(1-p)^{m_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}.$ (22)
Similarly, using the identity $(1-p)^{2G}=\left((1-p)^{2}\right)^{G}$, we obtain
$\mathbb{E}\big[(1-p)^{2G}\mid M_{0}=m_{0}\big]=\left(1-\big[1-(1-p)^{2}\big]\cdot\rho\right)^{N-{\color[rgb]{0,0,0}k}}=\left(1-(2p-p^{2})(1-p)^{m_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}.$ (23)
Taking expectations over $M_{0}$ (using (22) and (23)), we obtain
$\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]=\mathbb{E}\left[(T-M_{0})\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right],$
and
$\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]=\mathbb{E}\left[(T-M_{0})(T-M_{0}-1)\left(1-(2p-p^{2})(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right].$
To simplify these expressions, we now apply the following bounds. First, for all $M_{0}\geq 0$, since $(1-p)^{M_{0}}\in[0,1]$ and $p>0$, we have
$1-p(1-p)^{M_{0}}\geq 1-p.$
Then, since $x\mapsto x^{N-{\color[rgb]{0,0,0}k}}$ is monotonically increasing on $[0,1]$, it follows that
$\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\geq(1-p)^{N-{\color[rgb]{0,0,0}k}}.$
Therefore,
$\mathbb{E}\big[(T-M_{0})(1-p)^{G}\big]$ $=\mathbb{E}\left[(T-M_{0})\left(1-p(1-p)^{M_{0}}\right)^{N-{\color[rgb]{0,0,0}k}}\right]$
$\geq\mathbb{E}[T-M_{0}]\cdot(1-p)^{N-{\color[rgb]{0,0,0}k}}=T(1-q_{0})(1-p)^{N-{\color[rgb]{0,0,0}k}},$ (24)
For the second term, we simply upper-bound everything, such that, $(T-M_{0})(T-M_{0}-1)\leq T^{2}$ and $(1-p)^{2G}\leq 1$, so
$\mathbb{E}\big[(T-M_{0})(T-M_{0}-1)(1-p)^{2G}\big]\leq T^{2}.$ (25)
Substituting (24) and (25) into (B), we conclude the proof.
B-A Proof of (18)
Set $n\triangleq T-m_{0}$ and $\theta\triangleq a(1-p)^{g}$. Conditioned on $(M_{0},G)=(m_{0},g)$, there are exactly $n$ positive tests, and in each positive test, the event “test witnesses defective item ii” occurs with probability $\theta$, independently across tests. Therefore,
$L_{i}\,\big|\,(M_{0}=m_{0},G=g)\sim\mathrm{Bin}(n,\theta).$
By definition, the event $\{L_{i}=0\}$ means that none of the $n$ positive tests is a witness test for item $i$. Since these $n$ witness events are independent and each fails with probability $1-\theta$, we obtain
$P(L_{i}=0\mid M_{0}=m_{0},G=g)$
[…]
$\overset{(b)}{=}\prod_{t=1}^{n}P(\text{no witness in test $t$}\mid M_{0}=m_{0},G=g)$
$\overset{(c)}{=}\prod_{t=1}^{n}(1-\theta)$
$=(1-\theta)^{n}$
$=\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}},$
where (a) follows since this is the definition of $L_{i}=0$, (b) due to the independence across tests, and (c) as each test is a witness w.p. $\theta$. Finally, we have
$P(L_{i}>0\mid M_{0}=m_{0},G=g)=1-P(L_{i}=0\mid M_{0}=m_{0},G=g)=1-\bigl(1-a(1-p)^{g}\bigr)^{T-m_{0}},$
which yields (18).
B-B Binomial Bound
Lemma 2.
Let $u\in[0,1]$ and $n\in\mathbb{N}$. Then
$1-(1-u)^{n}\;\geq\;nu-\frac{n(n-1)}{2}u^{2}.$
Proof.
The function $f(u)=(1-u)^{n}$ is analytic on $[0,1]$, and has a Taylor expansion around $u=0$
$(1-u)^{n}=1-nu+\frac{n(n-1)}{2}u^{2}+R_{3}(u),$ (26)
where $R_{3}(u)$ denotes the third-order remainder in the Taylor expansion. For $n\geq 3$, the third derivative of $f$ is
$f^{(3)}(u)=-n(n-1)(n-2)(1-u)^{n-3}\leq 0\quad\text{for all }u\in[0,1].$
By the Lagrange form of the remainder, there exists $\xi\in(0,u)$ such that
$R_{3}(u)=\frac{f^{(3)}(\xi)}{3!}\,u^{3}\leq 0.$ (27)
For $n=1,2$, the expansion is exact and $R_{3}(u)\equiv 0$. Now, we multiply the identity in (26) by 1-1 and reorder terms to obtain
$1-(1-u)^{n}=nu-\frac{n(n-1)}{2}u^{2}-R_{3}(u).$
Since $R_{3}(u)\leq 0$ for all $u\in[0,1]$ as given in (27), we have $R_{3}(u)\geq 0$, and therefore
$1-(1-u)^{n}\geq nu-\frac{n(n-1)}{2}u^{2},$
which completes the proof. ∎
Appendix C Probability of Error (Proof of Lemma 1)
@@
Let $i\in\left[K\right],j\in\left[\binom{N-K}{i}\right$]. Let $\tilde{\mathcal{I}}^{j}_{dif}$ denote the set of all indices $\omega$ that result in a the set of FN of size $i$ as in $j$, i.e.
$\tilde{\mathcal{I}}^{j}_{dif}=\left\{\omega\in\left[\binom{N}{K}\right]:\quad FN\left(\omega,\omega^{*}\right)=\Omega_{i}\left(j\right)\right\}.$ (28)
@@
Recall $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. To visualize this definition, consider the example in Section IV-A. For $S_{\omega^{\star}}=\left\{1,2,3\right\}$, let $j$ be the index such that $\Omega_{2}\left(j\right)=\left\{2,3\right\}$. In this case, $\tilde{\mathcal{I}}^{j}_{dif}$ has all indices that point to $\left\{1,4,5\right\},\left\{1,4,6\right\},\left\{1,4,7\right\},\left\{1,5,6\right\},\left\{1,5,7\right\},\left\{1,6,7\right\}$.
@@
By substituting Eq. (30) in Eq. (C), it follows that
$\hskip-8.5359ptP\left(E_{i}^{j}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$ (31)
$\hskip-8.5359pt\leq\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)\frac{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}P_{W}\left(\omega^{\prime}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}P_{W}\left(\omega^{*}\right)^{s}}$
$\hskip-8.5359pt=\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}$
$\hskip-8.5359pt\overset{(a)}{\leq}\left[\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)}{P_{W}\left(\omega^{*}\right)^{s}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s}}\right]^{\rho}$
$\hskip-8.5359pt=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}},$
$\hskip-8.5359pt\hskip 18.49988pt\forall s>0,\rho\in\left[0,1\right]$
@@
$\hskip-8.5359ptP\left(E_{i}^{j}\lvert W=\omega^{*}\right)$ (32)
$\hskip-8.5359pt=\sum_{\mathbf{X}_{S_{\omega^{\star}}}}\sum_{Y^{T}}P\left(\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$\hskip-8.5359pt=\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(\mathbf{X}_{TP},Y^{T}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot P\left(E_{i}\lvert W=\omega^{*},\mathbf{X}_{S_{\omega^{\star}}},Y^{T}\right)$
$\hskip-8.5359pt\overset{(a)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{s}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{s\rho}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)\cdot P\left(\mathbf{X}_{FN}\right)\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{s}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)^{s\rho}}$
$\hskip-8.5359pt\overset{(b)}{\leq}\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right))}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\sum_{\mathbf{X}_{FN}}P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right)P\left(\mathbf{X}_{FN}\right)\cdot\frac{\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{\rho}}{P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FN}\right))^{\frac{\rho}{1+\rho}}}$
$\hskip-8.5359pt=\frac{\left(\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}\right)^{\rho}}{P_{W}\left(\omega^{*}\right)^{\frac{\rho}{\left(1+\rho\right)}}}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$\hskip-8.5359pt=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho}$
$\hskip-8.5359pt=\left[\frac{\sum_{\omega^{\prime}\in\tilde{\mathcal{I}}^{j}_{dif}}P_{W}\left(\omega^{\prime}\right)^{\frac{1}{\left(1+\rho\right)}}}{P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}}\right]^{\rho}\cdot\sum_{Y^{T}}\sum_{\mathbf{X}_{TP}}\left(\sum_{\mathbf{X}_{FP}}P\left(\mathbf{X}_{FP}\right)P\left(Y^{T},\mathbf{X}_{TP}\lvert\mathbf{X}_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},$
@@
$\begin{split}&\mathcal{T}\mathcal{P}_{i}=\left\{S_{1}:\lvert S_{1}\rvert=K-i\right\},\\ &\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)=\left\{S_{2}:\lvert S_{2}\rvert=i,\;S_{2}\cap S_{1}=\emptyset\right\},\\ &\hskip-36.98866pt\text{and}\\ &\omega\in\left[\binom{N}{K}\right]\;s.t.\;S_{\omega}=S_{1}\cup S_{2}.\end{split}$
@@
$\hskip-17.07182pt\binom{N}{K-i}\cdot\binom{N-\left(K-i\right)}{i}$